# Prot2Chat: Protein LLM with Early Fusion of Sequence and Structure

### v1

@@ 3.1 Overview of Prot2Chat
As illustrated in Figure 2, our model leverages the 3D coordinate structure information of protein residue atoms along with the protein sequence. The protein encoder processes this information to generate a protein embedding representation, which is then transformed into a text-aligned protein prompt via the protein-text adapter. This prompt, combined with the input question, serves as the input to the large language model (LLM), ultimately producing the output answer.
@@ 3.2 Sequence and Structure Fused Protein Encoder
[…]
@@
where Linear, Encoderlayers and Decoderlayers are all modules in ProteinMPNN, $E$ is the structual features and $n$ is the number of protein residues and $Dc$ is the embedding dimension of the protein encoder.
We find that a variant version of ProteinMPNN additionally contains the Embedding module for protein sequences $S$, and the Decoderlayer is sequence-aware, which changes Equation 4 to:
@@
To be specific, the Decoderlayer combines the node embeddings, the output features of the encoder, and the sequence embeddings to construct the context features for decoding through neighborhood feature aggregation. Using the message passing mechanism, the hidden state $h_{V}$ is continuously updated through multi-layer stacking to simulate the dynamic generation process of the decoder.
On this basis, we further initialized $h_{V}$ with the off-the-shelf sequence embedding to make the sequence and structural information early fused. Specifically, Equation 2 is replaced by:
@@
In particular, all of the model weights we used are from ProteinMPNN, with no need for additional training. The final protein node vector ${h}_{V}$ is used as the protein feature for the protein-text adapter. Inspired by InstructPLM Qiu et al. (2024), we concatenate the protein representations encoded by nine released ProteinMPNN models. As a result, $Dc=128*9=1152$.
3.3 Protein-Text Adapter
We implemented a protein-text adapter to semantically align the information obtained by the protein encoder with natural language, and provided both types of information as input to the large model simultaneously. The feature $h_{V}$ obtained from the modified ProteinMPNN model is output through the linear projection layer, positional encoding and the cross-attention mechanism. Specifically, the input protein feature first passes through a linear projection layer $W_{\text{proj}}$ to transform the input features to the target output dimension $D_{\text{o}}$:
@@ 3.3 Protein-Text Adapter
where $PE$ is the Dynamic Positional Encoding Vaswani (2017) used to capture the positional information of amino acids.
The complete protein encoding is too long. We adopt the idea of BLIP-2 to extract the important semantic features from it. In particular, we introduce $n_{\text{q}}$ learnable queries $\boldsymbol{Q}\in\mathbb{R}^{n_{\text{q}}\times D_{\text{o}}}$ and also apply $PE$ to them. Then the multi-head attention layer is adapted to capture key protein information based on queries:
@@
where $Q^{k}=\boldsymbol{Q}W_{Q}^{k}$, $K^{k}=X_{\text{proj}}W_{K}^{k}$ and $V^{k}=X_{\text{proj}}W_{V}^{k}$ stand for queries, keys and values. Among them, $W_{Q}^{k},W_{K}^{k},W_{V}^{k}$ are related parameters, $k$ is the head index. The multi-head attention outputs are then concatenated across heads and linearly transformed to the final protein prompt:
@@ 3.4 LLM Decoder
We combine $X_{\text{protein}}$ representing protein information with the text question and input it into the existing LLM to obtain the response.
@@
To improve domain adaptability, we fine-tuned the LLM with LoRA while training the adapter. The number of adapter training parameters is 89,702,400, while the number of LLaMA3 training parameters is 3,407,872. The total number of training parameters of the Prot2Chat is 93M. Meanwhile, the number of training parameters of BiomedGPT is 3B, that of ProtT3 is approximately 2B, the range of trainable parameters of Prot2Text is from 256M to 898M, and that of FAPM is 188M. This also shows the advantage of our model in significantly reducing the computational cost. Besides, we adpot the CrossEntropyLoss function as common.
@@ 4.1 Datasets
We evaluated the performance of our model using the Mol-Instructions Fang et al. (2023) and UniProtQA Luo et al. (2023) datasets. Mol-Instructions is a comprehensive instruction dataset specifically designed for the biomolecular domain, comprising both human-constructed and machine-generated data. For our experiments, we utilized the protein-oriented instruction subset, which is primarily derived from entries in the UniProtKB/Swiss-Prot database UniProt Consortium (2018). This subset encompasses tasks such as predicting protein domains, functions, and activities. Our model was predominantly trained on this dataset.
Additionally, we employed a portion of the UniProtQA dataset, introduced by BiomedGPT Luo et al. (2023), to assess the generalization capability of our model through zero-shot evaluation. Fine-tuning was also performed on this dataset to further validate performance. UniProtQA consists of textual descriptions of proteins, including their functions and properties. It was curated by UniProt UniProt Consortium (2018) and includes four types of questions related to protein function, formal name, protein family, and subcellular location. The detailed sizes of our dataset splits are provided in Table 2.
To incorporate structural and sequence information, we retrieved the corresponding PDB files for proteins listed in Swiss-Prot. These PDB files, combined with the associated questions, serve as the input to Prot2Chat.
@@ 4.2 Baselines
  • LLaMA3: LLaMA3 Dubey et al. (2024) series has attracted a lot of research attention due to its outstanding capabilities in the general domain. We use LLaMA3-8B-Instruct and input the protein sequence and question to achieve zero-shot.
@@
  • BioMedGPT-10B: BioMedGPT is a domain-specific LLM trained on a large selection of corpora of human scientific knowledge. The model encodes the protein sequence with the ESM-3B Lin et al. (2022) and uses BioMedGPT-LM-7B as the decoder to generate the response.
@@
In order to make a fair comparison, we used the same input template for each model through prompt engineering, so that the model could output the answer as we instructed. Some other open-source models in related fields, such as FAPM, Prot2Text, and InstructProtein, were not included in the comparison because these models complete protein Q&A tasks based on classification, or there are differences in the formats of the model’s inputs and outputs, as shown in Table 1. Therefore, it is hard to conduct comparisons.
@@ 4.3 Model Setting
We jointly trained the protein adapter and LLaMA3 on the Mol-Instructions dataset, employing full training for the adapter and LoRA (Low-Rank Adaptation) fine-tuning for LLaMA3. Inspired by BLIP-2 Li et al. (2023), we configured the adapter with 256 queries. For the LoRA setup, we set the rank $r$ to 8, LoRA alpha to 16, and targeted the ”q_proj” and ”v_proj” modules, with a LoRA dropout rate of 0.1. The adapter comprises 89,702,400 trainable parameters, while LLaMA3 has 3,407,872 trainable parameters.
We utilized the Adam optimizer and implemented gradient accumulation to optimize the training process. The initial learning rate was set to $10^{-4}$, with a batch size of 2 and a maximum context length of 1024 tokens, which includes both the question and answer text. The model was fine-tuned for 2 epochs, with each training session requiring approximately 1600 hours on an NVIDIA RTX 3090 GPU.
@@ 4.4 Evaluation Metrics
To evaluate the effectiveness of the model’s text generation, we employed performance metrics including BLEU Papineni et al. (2002) and ROUGE Lin (2004). Additionally, we utilized the prompt templates shown in Table 3 to simultaneously input the target text and the text generated by different models into the online KIMI model. This allowed us to determine which generated answer was closer to the target text, thereby assisting in the evaluation of the model’s output quality. The results, as shown in Table 6, demonstrate that our model outperforms other comparative models overall. Furthermore, we conducted expert manual evaluations, where professional biology PhDs ranked the responses of different models based on their alignment with the target text. The results of these evaluations (Table 7) are consistent with those obtained using KIMI.
@@ 4.5 Main Results
The results of the assessment are shown in Table 4 and Table 5. The results show that there is a significant modal gap between the protein sequence and the natural language, and it is incomprehensible to directly input the protein sequence as text and problem into the large language model, resulting in the disordered and meaningless response of the model. Aligning protein language with human language is an effective solution to solve this problem, and we have observed that the performance of the model trained with the addition of structural information is significantly better than that of other models. This suggests that we can directly use the structure and sequence information of proteins to help language models better understand and generate technical terms through feature space alignment. Establishing a link between proteins and natural language can provide guidance for researchers in their research on unknown proteins. The results of our model’s prediction on the UniprotQA Luo et al. (2023) dataset verify the generalization performance of our model. Further illustrating the role of our model in protein Q&A tasks.
@@ 4.6 Ablation Experiments
As shown in Table 4, we conducted a comparative analysis to evaluate the impact of incorporating protein structure during training, thereby highlighting the importance of structural information for protein understanding. We also examined the effect of fine-tuning LLMs, revealing that training adapters alone was insufficient for achieving optimal performance. Furthermore, by comparing the results obtained from training with sequence text alone versus training with both sequence and structural information, we observed that LLMs struggle to effectively interpret protein sequence information presented in textual form. These findings underscore the necessity of integrating protein structural information alongside a unified multimodal alignment module, which enables LLMs to better understand and generate domain-specific technical terms.
@@ 4.7 Case Study
As shown in Table 8, the results we obtained using the protein Q&A test dataset indicate that the model trained with the addition of protein structure information generates responses significantly better than the model trained with protein sequence. Further prove that it is difficult for LLM to discover the corresponding protein functions and effects only through the text information of protein sequences, and thus draw the conclusion that the protein structure is vital for the comprehensive understanding of proteins.
@@ 5 Conclusion
This paper introduces a novel protein Q&A model that addresses the limitations of traditional protein Q&A tasks by leveraging the modified ProteinMPNN and an adapter module to integrate protein structure and sequence information with natural language text. By harnessing the flexible generation capabilities of large language models (LLMs), our approach effectively bridges the gap between protein data and textual understanding. Experimental results demonstrate that the model trained with structural information significantly outperforms other baseline models, underscoring the critical role of protein structure in protein understanding and analysis. Furthermore, the results validate the effectiveness and strong generalization ability of our model in protein Q&A tasks.

### v2

@@ 3.1 Overview of Prot2Chat
As illustrated in Figure 2, the overall architecture of our model comprises three components: the protein encoder, the text-aware protein-text adapter, and the LLM. Our model leverages the 3D coordinate structure information of protein residue atoms along with the protein sequence and natural language question text. The protein encoder processes protein information to generate a protein embedding representation, which is then transformed into a text-aligned protein prompt via the text-aware protein-text adapter, with question text information being fused simultaneously. This prompt, combined with the input question, serves as the input to the large language model, ultimately producing the output answer.
@@ 3.2 Sequence and Structure Fused Protein Encoder
We modified the structure encoder ProteinMPNN to realize the sequence and structure early-fusion. ProteinMPNN is originally dedicated to design protein sequences based on backbone structures. The input is structural information $E$, namely the 3D coordinates of protein residue atoms ($N$: amino nitrogen, $C\alpha$: alpha carbon, $C$: carbonyl carbon and $O$: carbonyl oxygen). Then protein features $h_{V}$ are obtained through the encoder layer and the decoder layer.
@@
where Linear, Encoderlayers and Decoderlayers are all modules in ProteinMPNN, $E$ is the structual features and $n$ is the number of protein residues and $Dc$ is the embedding dimension of the protein encoder. We find that a variant version of ProteinMPNN additionally contains the Embedding module for protein sequences $S$, and the Decoderlayer is sequence-aware, which changes Equation 4 to:
@@
To be specific, the Decoderlayer combines the node embeddings, the output features of the encoder, and the sequence embeddings to construct the context features for decoding through neighborhood feature aggregation. Using the message passing mechanism, the hidden state $h_{V}$ is continuously updated through multilayer stacking to simulate the dynamic generation process of the decoder.
On this basis, we further initialized $h_{V}$ with the off-the-shelf sequence embedding to make the sequence and structural information early-fused. Specifically, Equation 2 is replaced by:
@@
In particular, all of the model weights we used are from ProteinMPNN, with no need for additional training. The final protein node vector ${h}_{V}$ is used as the protein feature embedding for the text-aware protein-text adapter. Inspired by InstructPLM qiu2024instructplm , we concatenate the protein representations encoded by nine released ProteinMPNN models. As a result, $Dc=128*9=1152$.
3.3 Text-Aware Protein-Text Adapter
We implemented a text-aware protein-text adapter to semantically align the information obtained from the protein encoder with natural language, incorporating question text information simultaneously. Subsequently, this fused information is provided as a soft prompt to the LLM. The protein embedding $h_{V}$ obtained from the modified ProteinMPNN model is output through the linear projection layer, positional encoding and the cross-attention mechanism. During the above-mentioned process, text information is fused simultaneously, as shown in Figure  3.
Figure 3: Detail of the Text-Aware Protein-Text Adapter. The adapter takes protein embedding and question vector as inputs. It integrates text information at an early stage through a set of learnable queries. Subsequently, using these queries, the protein embedding is used as keys/values for cross-attention calculations. This cross-attention module compresses the protein information into a fixed length and captures the key protein information according to the queries infused with question text information, resulting in soft prompt to assist the LLM in generating more accurate responses.
To be specific, the input protein feature first passes through a linear projection layer $W_{\text{proj}}$, the aim is to transform the input features to the target output dimension $D_{\text{o}}$:
@@ 3.3 Text-Aware Protein-Text Adapter
where $PE$ is the Dynamic Positional Encoding vaswani2017attention used to capture the positional information of amino acids.
Meanwhile, we utilize the question text and the LLM to obtain the hidden state $Qh_{t}$ of the last text token as the question vector, so as to achieve prompt compression and carry out the early-fusion of text and protein information, enabling the model to perceive question text information earlier:
$Qh_{\text{t}}=\text{LLM}(\text{question}).hiddenstate$ (8)
The complete protein encoding is too long. We adopt the idea of BLIP-2 to extract the important semantic features from it. In particular, we introduce $n_{\text{q}}$ learnable queries $\boldsymbol{Q}\in\mathbb{R}^{n_{\text{q}}\times D_{\text{o}}}$, incorporate the question vector $Qh_{t}$ obtained from the question text, and also apply the positional encoding (PE) to them.
Then the multi-head cross-attention layer is adapted to capture key protein information based on queries:
@@
where $Q^{k}=\boldsymbol{Q}W_{Q}^{k}$, $K^{k}=X_{\text{proj}}W_{K}^{k}$ and $V^{k}=X_{\text{proj}}W_{V}^{k}$ stand for queries, keys and values. Among them, $W_{Q}^{k},W_{K}^{k},W_{V}^{k}$ are related parameters, $k$ is the head index.
Then, the outputs of multi-head cross-attention are concatenated across heads and linearly transformed by ${W}_{out}$ into the final protein prompt integrated with text information, which serves as the soft prompt for the LLM, where $M$ is the number of attention heads:
@@ 3.4 LLM Decoder
We combine $X_{\text{protein}}$, which integrates text and protein information, with the text question, and then input this combination into the existing LLM to obtain a response.
@@
To improve domain adaptability, we fine-tuned the LLM with LoRA while training the adapter. The number of adapter training parameters is 106,483,712, while the number of LLaMA3 training parameters is 3,407,872.
The total number of training parameters of the Prot2Chat is 109M. Meanwhile, the number of training parameters of BiomedGPT-10B is 3B, that of Evola-10B is approximately 1.7B, the range of trainable parameters of Prot2Text is from 256M to 898M, and that of FAPM is 188M. This also shows the advantage of our model in significantly reducing the computational cost. Besides, we adpot the CrossEntropyLoss function as common.
@@ 4.1 Datasets
We evaluated the performance of our model using the Mol-Instructions fang2023mol and UniProtQA luo2023biomedgpt datasets. For our experiments, we utilized the protein-oriented instruction subset in Mol-Instructions, which is primarily derived from entries in the UniProtKB/Swiss-Prot database uniprot2018uniprot . This subset encompasses tasks such as predicting protein domains, functions, and activities. Our model was predominantly trained on this dataset.
Additionally, we employed a portion of the UniProtQA dataset, introduced by BiomedGPT luo2023biomedgpt , to assess the generalization capability of our model through zero-shot evaluation. Fine-tuning was also performed on this dataset to further validate performance. UniProtQA consists of textual descriptions of proteins, including their functions and properties. The detailed sizes of our dataset splits are provided in Table 2. It is worth noting that there are significant differences in the text style and length between the UniProtQA and Mol-Instructions datasets. This directly leads to large differences in the traditional metrics of the subsequent evaluation test results. To incorporate structural and sequence information, we retrieved the corresponding PDB files for proteins listed in Swiss-Prot. These PDB files, combined with the associated questions, serve as the input to Prot2Chat.
@@ 4.2 Baselines
  • LLaMA3: LLaMA3 dubey2024llama series has attracted a lot of research attention due to its outstanding capabilities in the general domain. We use LLaMA3-8B-Instruct and input the protein sequence and question to achieve zero-shot.
@@
  • BioMedGPT-LM-10B: BioMedGPT luo2023biomedgpt is a domain-specific LLM trained on a large selection of corpora of human scientific knowledge. The model encodes the protein sequence with the ESM-3B lin2022language and uses BioMedGPT-LM-7B as the decoder to generate the response.
  • Evola-10B: Evola zhou2025decoding is trained on hundreds of millions of protein Q&A pairs and an AI-generated dataset of 150 billion words. It adopts Retrieval-Augmented Generation (RAG) to integrate external knowledge and performs excellently in generating accurate, detailed, and context-relevant answers about protein functions, providing expert-level insights for proteomics and functional genomics research. Here, we use its model with 10B parameters as the comparative model.
@@
In order to make a fair comparison, we used the same input template for each model through prompt engineering, so that the model could output the answer as we instructed. Some other open-source models in related fields, such as FAPM, Prot2Text, and InstructProtein, were not included in the comparison because these models complete protein Q&A tasks based on classification, or there are differences in the formats of the model’s inputs and outputs, as shown in Table 1. Therefore, it is hard to conduct comparisons. In addition, the ProtChatGPT model was not included in the comparison because accessible resources for this model could not be found.
@@ 4.3 Model Setting
We jointly trained the text-aware protein-text adapter and LLaMA3 on the Mol-Instructions dataset, employing full training for the adapter and LoRA fine-tuning for LLaMA3. Inspired by InstructPLM qiu2024instructplm , we configured the adapter with 256 queries and integrate text information on each query. For the LoRA setup, we set the rank $r$ to 8, LoRA alpha to 16, and targeted the ”q_proj” and ”v_proj” modules, with a LoRA dropout rate of 0.1. The adapter comprises 106,483,712 trainable parameters, while LLaMA3 has 3,407,872 trainable parameters. We utilized the Adam optimizer and implemented gradient accumulation to optimize the training process. The initial learning rate was set to $10^{-4}$, with a batch size of 2 and a maximum context length of 1024 tokens, which includes both the question and answer text. The model was fine-tuned for 2 epochs, with each training session requiring approximately 1600 hours on an NVIDIA RTX 3090 GPU.
@@ 4.4 Evaluation Metrics
To evaluate the effectiveness of the model’s text generation, we employed performance metrics including BLEU papineni-etal-2002-bleu and ROUGE lin-2004-rouge . Additionally, the target text and the texts generated by different models are simultaneously input into the online KIMI model, using the following prompts: Please select the sentence that is closest and most accurate in meaning to the Target sentence and rank sentences A, B, C and D accordingly, with the first place meaning the most identical and accurate in meaning to the Target. Only provide the answer, for example, ‘B D A C’ or ‘A C D B’, etc. Target: {Ground Truth} A: {Model1 Generated Text} B: {Model2 Generated Text} C: {Model3 Generated Text} D: {Model4 Generated Text} This allowed us to determine which generated answer was closer to the target text, thereby assisting in the evaluation of the model’s output quality. Furthermore, we conducted expert manual evaluations, where professional biology PhDs ranked the responses of different models based on their alignment with the target text.
@@ 4.5 Main Results
The results of the assessment are shown in Table 3 to Table 7. The BLEU-2 and ROUGE-L scores of Prot2Chat exceed those of models such as BiomedGPT and Evola. The results show that there is a significant modal gap between the protein sequence and the natural language, and it is incomprehensible to directly input the protein sequence as text and problem into the LLM, resulting in the disordered and meaningless response of the model. Aligning protein language with human language is an effective way to solve this problem. Moreover, we have observed that the performance of the model trained through the early-fusion of structural information and sequences is significantly superior to that of other models leveraging either sequences or structures. This suggests that we can directly use the structure and sequence information of proteins to help language models better understand and generate technical terms through feature space alignment. In addition, the early-fusion of text information and protein information also demonstrates a significant effect. Unification of multimodal protein information and establishment of a connection between proteins and natural language can provide guidance for researchers in their study of unknown proteins. The results of our model’s prediction on the UniProtQA luo2023biomedgpt dataset verify the generalization performance of our model. On the other hand, the evaluation results of online large language models, as shown in Table 6, demonstrate that our model outperforms other comparative models overall and the results of expert manual evaluations (Table 7) are consistent with those obtained using KIMI. Further illustrating the role of our model in protein Q&A tasks.
@@ 4.6 Ablation Experiments
As shown in Table 4, we carried out a comparative analysis to evaluate the impact of incorporating protein structure information during the training process. Comparing with the results of the model trained with only the sequence further highlights the importance of structural information for understanding proteins. We also studied the effect of fine-tuning LLMs. The results show that training only the adapter is insufficient to achieve optimal performance. By comparing the results obtained from training in the cases of early-fusion and non-early-fusion of text and protein information, we found that the early-fusion of information plays a non-negligible role in the understanding of proteins and the generation of large language models, and its effect is significantly better than that of late-fusion. These findings emphasize the necessity of integrating protein structure information and a unified multimodal alignment module, which enables large language models to better understand and generate domain-specific technical terms.
@@ 4.7 Case Study
As shown in Table 8, the results obtained from our use of the protein Q&A test dataset indicate that the responses generated by the model trained by adding protein structure and early-fusing text information are significantly superior to those of other trained models (taking LLaMA3-FT as an example here). This further demonstrates that protein structure is of crucial importance for the comprehensive understanding of proteins, the early-fusion and alignment of multimodal information plays a non-negligible role in the understanding of proteins and the generation of large language models.
@@ 5 Conclusion
This paper introduces a novel protein Q&A model that addresses the limitations of traditional protein Q&A tasks by leveraging the modified ProteinMPNN and a text-aware adapter module to integrate protein structure and sequence information with natural language text. By harnessing the flexible generation capabilities of large language models (LLMs), our approach effectively bridges the gap between protein data and textual understanding. Experimental results demonstrate that the model trained by adding protein structure and early-fusing text information significantly outperforms other baseline models. This not only highlights the crucial role of protein structure in protein understanding and analysis but also proves the effectiveness of the early-fusion of protein information and text. Furthermore, the results validate the effectiveness and generalization ability of our model in protein Q&A tasks.